OpenStudy (anonymous):

Y-5=4(x+2) in standard form

OpenStudy (anonymous):

y-5=4x+8 4x-y=-5-8 4x-y=-13

OpenStudy (anonymous):

then move the y over and the 13 over and get y=4x+13